The following circuit operates if and only if there is a path of functional devices from left to right. Assume devices fail independently and that the probability of failure of each device is as shown. What is the probability that the circuit operates? 4. 0.010.01 0.1
Answer #1

The probability that the circuit operates

= P(upper left device is working, upper right device is working, lower device is working) + P(upper left device is not working, upper right device is not working, lower device is working)

= [(1 - 0.01)*(1 - 0.01)*(1-0.1)] + [0.01*0.01*(1-0.1)]

=(0.99*0.99*0.9) + ( 0.01*0.01*0.9)

= 0.88209 + 0.00009

= 0.88218 (ans)